# Dual Ascent Diffusion for Inverse Problems

### v1

@@ A.1 Statistical Significance
FFHQ ImageNet Task Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) Super Resolution 4× DDiff (ours) 30.07$\pm$0.41 0.824$\pm$0.008 0.211$\pm$0.009 (2.85±$\pm 0.06$)$\cdot 10^{-3}$ 25.81$\pm$0.72 0.656$\pm$0.029 0.396$\pm$0.029 (3.83±$\pm 0.29$)$\cdot 10^{-3}$ DAPS 29.34$\pm$0.33 0.783$\pm$0.006 0.190$\pm$0.006 (2.87±$\pm 0.06$)$\cdot 10^{-3}$ 25.44$\pm$0.56 0.636$\pm$0.017 0.295$\pm$0.011 (2.87±$\pm 0.36$)$\cdot 10^{-1}$ Inpainting (Box) DDiff (ours) 24.88$\pm$0.50 0.831$\pm$0.005 0.110$\pm$0.005 (7.75±$\pm 9.21$)$\cdot 10^{-3}$ 21.15$\pm$0.68 0.743$\pm$0.009 0.240$\pm$0.011 (1.19±$\pm 0.29$)$\cdot 10^{-2}$ DAPS 24.52$\pm$0.40 0.742$\pm$0.006 0.174$\pm$0.006 (9.86±$\pm 6.76$)$\cdot 10^{-3}$ 21.22$\pm$0.70 0.714$\pm$0.007 0.230$\pm$0.011 (2.05±$\pm 0.24$)$\cdot 10^{-1}$ Inpainting (Random) DDiff (ours) 33.08$\pm$0.37 0.877$\pm$0.004 0.050$\pm$0.003 (2.05±$\pm 0.15$)$\cdot 10^{-2}$ 28.39$\pm$0.76 0.758$\pm$0.019 0.136$\pm$0.018 (2.41±$\pm 0.35$)$\cdot 10^{-2}$ DAPS 30.76$\pm$0.27 0.801$\pm$0.005 0.156$\pm$0.003 (2.93±$\pm 0.16$)$\cdot 10^{-2}$ 27.32$\pm$0.63 0.725$\pm$0.013 0.189$\pm$0.010 (7.88±$\pm 0.70$)$\cdot 10^{-2}$ Gaussian Deblurring DDiff (ours) 28.87$\pm$0.43 0.800$\pm$0.010 0.119$\pm$0.005 (2.60±$\pm 0.07$)$\cdot 10^{-3}$ 22.29$\pm$0.84 0.471$\pm$0.039 0.415$\pm$0.035 (4.68±$\pm 0.39$)$\cdot 10^{-3}$ DAPS 29.63$\pm$0.36 0.789$\pm$0.006 0.177$\pm$0.005 (2.67±$\pm 0.06$)$\cdot 10^{-3}$ 25.90$\pm$0.64 0.658$\pm$0.020 0.269$\pm$0.011 (2.37±$\pm 0.31$)$\cdot 10^{-1}$ Motion Deblurring DDiff (ours) 28.24$\pm$0.38 0.785$\pm$0.009 0.129$\pm$0.005 (5.82±$\pm 0.27$)$\cdot 10^{-3}$ 24.16$\pm$0.65 0.585$\pm$0.027 0.242$\pm$0.014 (7.97±$\pm 0.60$)$\cdot 10^{-3}$ DAPS 29.17$\pm$0.36 0.797$\pm$0.007 0.186$\pm$0.005 (5.91±$\pm 0.28$)$\cdot 10^{-3}$ 26.61$\pm$0.65 0.710$\pm$0.020 0.241$\pm$0.011 (1.82±$\pm 0.25$)$\cdot 10^{-1}$ Phase Retrieval DDiff (ours) 29.94$\pm$0.88 0.816$\pm$0.025 0.120$\pm$0.019 1.469$\pm$0.002 18.54$\pm$1.23 0.494$\pm$0.045 0.262$\pm$0.029 1.465$\pm$0.002 DAPS 29.60$\pm$0.88 0.768$\pm$0.020 0.182$\pm$0.018 1.477$\pm$0.002 20.23$\pm$1.27 0.449$\pm$0.044 0.397$\pm$0.032 1.469$\pm$0.002 Nonlinear Deblurring DDiff (ours) 31.48$\pm$0.29 0.873$\pm$0.005 0.120$\pm$0.006 (2.74±$\pm 0.07$)$\cdot 10^{-3}$ 29.68$\pm$0.62 0.805$\pm$0.016 0.207$\pm$0.019 (3.55±$\pm 0.18$)$\cdot 10^{-3}$ DAPS 28.45$\pm$0.38 0.764$\pm$0.007 0.188$\pm$0.006 (4.16±$\pm 0.43$)$\cdot 10^{-3}$ 27.28$\pm$0.62 0.718$\pm$0.017 0.213$\pm$0.011 (2.64±$\pm 0.33$)$\cdot 10^{-1}$ HDR DDiff (ours) 26.05$\pm$0.68 0.873$\pm$0.007 0.129$\pm$0.009 (4.59$\pm$0.27)$\cdot 10^{-2}$ 26.50$\pm$0.56 0.800$\pm$0.018 0.108$\pm$0.014 (5.41$\pm$0.27)$\cdot 10^{-2}$ DAPS 27.39$\pm$0.56 0.846$\pm$0.010 0.163$\pm$0.012 (5.05±$\pm 0.18$)$\cdot 10^{-2}$ 26.10$\pm$0.79 0.825$\pm$0.021 0.171$\pm$0.019 (7.17±$\pm 0.75$)$\cdot 10^{-1}$
A.2 Relating DDiff and DiffPIR
@@
Here we show the equivalence of DDiff and DiffPIR (see Algorithm S1) under specific parameter choices. DiffPIR introduces $\zeta$ as a hyperparameter controlling the stochasticity of the algorithm and $\bar{\sigma}_{t}$ as a parameter controlling the strength of the likelihood step. Our algorithm is equivalent to the linearized version of DiffPIR when:
@@
A.3 DDiff Variants Algorithms
This section presents the precise algorithms corresponding to the three variants of DDiff discussed in the ablation study. Note that DDiff-HQS S4 is equivalent to DiffPIR S1 if we satisfy (i.-iii.) from section A.2.
Algorithm S2 Diff-PnP-HQS (no noise, no $\mathbf{u}$)
@@ A.3 DDiff Variants Algorithms
Algorithm S3 Diff-PnP-ADMM (no noise, with $\mathbf{u}$)
@@
Algorithm S4 DDiff-HQS (with noise, no $\mathbf{u}$)
@@
A.4 Inverse Problems Setup
@@
  • Super resolution: 4×4$\times$4 × downsampling factor
@@
  • HDR: 2×2$\times$2 × dynamic range
A.5 Hyperparameter Choices
@@
$\gamma_{t}=\gamma_{0}\cdot f(t_{\gamma})$ where $f(t_{\gamma})=\begin{cases}3.3&\text{for }t>t_{\gamma},\\ 0.1&\text{for }t\leq t_{\gamma}.\end{cases}$
To enhance sample quality, the parameter $\gamma_{t}$ is reduced at time step $t_{\gamma}$. This adjustment encourages the traversal to relax toward the prior manifold while diverging from the likelihood manifold. As a result, the generated samples exhibit reduced noise and improved visual fidelity.
@@
  • $\sigma_{t}=\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{1-\alpha_{t}/\alpha_{t-1}}$ (this choice allows the generative process to become a DDPM) was used for Gaussian deblurring, motion deblurring, nonlinear deblurring, and HDR.
@@
We report the hyperparameters used in the main experiments in Table S2 and S3.
Table S2: DDiff hyperparameter settings for different tasks on FFHQ.
Super Res. 4×4$\times$4 × Inpaint (Box) Inpaint (Rand.) Gaussian deblur
$\gamma_{0}$ 18 30 50 2.9
$t_{\gamma}$ 90 90 90 90
$t_{0}$ 1 1 1 50
Motion deblur Phase retrieval Nonlinear deblur HDR
$\gamma_{0}$ 2.9 38 2.5 3.5
$t_{\gamma}$ 90 90 90 90
$t_{0}$ 80 1 120 120
Table S3: DDiff hyperparameter settings for different tasks on ImageNet.
Super Res. 4×4$\times$4 × Inpaint (Box) Inpaint (Rand.) Gaussian deblur
$\gamma_{0}$ 18 50 50 1.8
$t_{\gamma}$ 90 500 90 90
$t_{0}$ 1 1 1 50
Motion deblur Phase retrieval Nonlinear deblur HDR
$\gamma_{0}$ 1.5 38 2.5 3.8
$t_{\gamma}$ 90 90 90 90
$t_{0}$ 80 1 120 100
A.6 More Qualitative Results
We provide additional qualitative results in Fig. S1 to S8.
Figure S1: Visual comparison of DDiff and baselines on super resolution task.
Figure S2: Visual comparison of DDiff and baselines on box inpainting task.
Figure S3: Visual comparison of DDiff and baselines on random inpainting task.
Figure S4: Visual comparison of DDiff and baselines on Gaussian deblurring task.
Figure S5: Visual comparison of DDiff and baselines on motion deblurring task.
Figure S6: Visual comparison of DDiff and baselines on phase retrieval task.
Figure S7: Visual comparison of DDiff and baselines on nonlinear deblurring task.
Figure S8: Visual comparison of DDiff and baselines on HDR task.

### v2

@@ Appendix C DDiff with Latent Diffusion Models
FFHQ ImageNet Task Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) Residual ($\downarrow$) Super Resolution 4× LatentDDiff 26.67 0.688 0.324 0.0040 28.65 0.664 0.411 0.0043 LatentDAPS 27.35 0.800 0.185 0.0038 25.01 0.669 0.284 0.0058 PSLD 24.24 0.639 0.289 0.0060 25.40 0.689 0.362 0.0056 ReSample 23.13 0.588 0.401 0.0065 22.57 0.567 0.382 0.0065 Inpainting (Box) LatentDDiff 20.83 0.617 0.208 0.0192 19.10 0.551 0.337 0.0162 LatentDAPS 23.86 0.799 0.192 0.0146 17.15 0.645 0.341 0.0372 PSLD 24.15 0.800 0.163 0.0095 20.08 0.687 0.479 0.0390 ReSample 19.56 0.621 0.190 0.0131 18.13 0.625 0.270 0.0214 Inpainting (Random) LatentDDiff 26.81 0.643 0.239 0.0035 29.15 0.665 0.341 0.0038 LatentDAPS 30.52 0.804 0.153 0.0033 27.31 0.769 0.177 0.0058 PSLD 30.14 0.776 0.228 0.0037 30.23 0.760 0.340 0.0042 ReSample 29.24 0.730 0.145 0.0036 27.03 0.721 0.155 0.0095 Gaussian Deblurring LatentDDiff 27.56 0.725 0.252 0.0027 26.06 0.850 0.310 0.0027 LatentDAPS 27.58 0.753 0.248 0.0026 24.67 0.637 0.356 0.0067 PSLD 22.98 0.600 0.331 0.0064 25.03 0.656 0.411 0.0058 ReSample 25.56 0.697 0.278 0.0045 25.45 0.683 0.267 0.0049 Motion Deblurring LatentDDiff 26.11 0.593 0.286 0.0028 25.54 0.637 0.330 0.0038 LatentDAPS 26.88 0.799 0.299 0.0027 26.34 0.721 0.304 0.0034 PSLD 21.81 0.642 0.357 0.0036 20.23 0.566 0.525 0.0067 ReSample 27.21 0.801 0.214 0.0025 26.65 0.706 0.244 0.0030 Nonlinear Deblurring LatentDDiff 27.06 0.683 0.234 0.0050 26.80 0.711 0.265 0.0041 LatentDAPS 27.87 0.695 0.250 0.0056 25.04 0.612 0.346 0.0057 ReSample 28.04 0.722 0.199 0.0052 25.89 0.621 0.245 0.0044 High Dynamic Range LatentDDiff 22.68 0.648 0.204 0.0145 21.69 0.662 0.201 0.0360 LatentDAPS 25.66 0.737 0.255 0.0154 23.32 0.589 0.280 0.0366 ReSample 25.21 0.702 0.197 0.0161 24.97 0.616 0.208 0.0381
Appendix D Statistical Significance
Building on the main results table, we report 95% confidence intervals for DDiff and DAPS in Table 4, highlighting the statistical significance of our improvement, compared to DAPS [43]. These statistics also serve as empirical evidence for the practical uniqueness of the fixed point established in Theorem 1; the narrow confidence intervals across all tasks indicate that independent initializations $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ yield tightly concentrated reconstructions.
Table 4: Quantitative evaluation with confidence intervals. We show the average metric values over 100 validation images and the corresponding 95% confidence intervals for DDiff and DAPS.
[…]
Appendix E Relating DDiff and DiffPIR
@@
Here we show the equivalence of DDiff and DiffPIR (see Algorithm 3) under specific parameter choices. DiffPIR introduces $\zeta$ as a hyperparameter controlling the stochasticity of the algorithm and $\bar{\sigma}_{t}$ as a parameter controlling the strength of the likelihood step. Our algorithm is equivalent to the linearized version of DiffPIR when:
@@
Appendix F DDiff Variants Algorithms
This section presents the precise algorithms corresponding to the three variants of DDiff discussed in the ablation study. Note that DDiff-HQS 6 is equivalent to DiffPIR 3 if we satisfy (i.-iii.) from Appendix E.
Algorithm 4 Diff-PnP-HQS (no noise, no $\mathbf{u}$)
@@ Appendix F DDiff Variants Algorithms
Algorithm 5 Diff-PnP-ADMM (no noise, with $\mathbf{u}$)
@@
Algorithm 6 DDiff-HQS (with noise, no $\mathbf{u}$)
@@
Appendix G Inverse Problems Setup
@@
  • Super resolution: $4\times$ downsampling factor
@@
  • HDR: $2\times$ dynamic range
Appendix H Hyperparameter Choices
@@
$\gamma_{t}=\gamma_{0}\cdot f(t_{\gamma})$ where $f(t_{\gamma})=\begin{cases}a&\text{for }t>t_{\gamma},\\ b&\text{for }t\leq t_{\gamma}.\end{cases}$
To enhance sample quality, the parameter $\gamma_{t}$ is reduced at time step $t_{\gamma}$. This adjustment encourages the traversal to relax toward the prior manifold while diverging from the likelihood manifold. As a result, the generated samples exhibit reduced noise and improved visual fidelity. For all pixel-space diffusion tasks, $a=3.3$ and $b=0.1$ worked well empirically. For latent-space diffusion tasks, we report the exact values in Table 5 and 6.
@@
  • $\sigma_{t}=\sqrt{(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}\sqrt{1-\bar{\alpha}_{t}/\bar{\alpha}_{t-1}}$ (this choice allows the generative process to become a DDPM) was used for Gaussian deblurring, motion deblurring, nonlinear deblurring, and HDR.
@@
The time threshold $t_{0}$ controls the reverse diffusion step to switch from stochastic to deterministic.
We report the hyperparameters used in the main experiments in Table 5 and 6.
Table 5: DDiff hyperparameter settings for different tasks on FFHQ.
Algorithm Task Super Res. $4\times$ Inpaint (Box) Inpaint (Rand.) Gaussian Deblur Motion Deblur Phase Retrieval Nonlinear Deblur HDR
DDiff $\gamma_{0}$ 18 30 50 2.9 2.9 38 2.5 3.5
$t_{\gamma}$ 90 90 90 90 90 90 90 90
$t_{0}$ 1 1 1 50 80 1 120 120
LatentDDiff $\gamma_{0}$ 100 200 400 60 60 - 500 150
$t_{\gamma}$ 90 600 200 90 150 - 90 600
$t_{0}$ 500 1 800 200 200 - 200 1
$a$ 3.3 25 30 3.3 3.3 - 3.3 30
$b$ 0.1 0.15 0.1 0.1 0.1 - 0.1 0.15
Table 6: DDiff hyperparameter settings for different tasks on ImageNet.
Algorithm Task Super Res. $4\times$ Inpaint (Box) Inpaint (Rand.) Gaussian Deblur Motion Deblur Phase Retrieval Nonlinear Deblur HDR
DDiff $\gamma_{0}$ 18 50 50 1.8 1.5 38 2.5 3.8
$t_{\gamma}$ 90 500 90 90 90 90 90 90
$t_{0}$ 1 1 1 50 80 1 120 100
LatentDDiff $\gamma_{0}$ 150 400 400 150 100 - 100 250
$t_{\gamma}$ 600 600 600 600 90 - 90 600
$t_{0}$ 1 1 1 1 200 - 200 200
$a$ 50 30 50 50 3.3 - 3.3 25
$b$ 0.15 0.1 0.15 0.15 0.1 - 0.1 0.15
Appendix I Baseline Details
We follow the baseline configurations reported in DAPS [43] for all methods we compare against: DPS [7], DDRM [21], DCDP [24], DiffPIR [46], RED-diff [27], PSLD [30], and ReSample [32]. For DAPS and DMPlug [40], we use the authors’ official implementations with default hyperparameters. All baselines use the same pretrained models (as specified in Sec. 4) and forward operators (Sec. G).
In our experiments, we use DDiff-1k for all inverse problems. In contrast, DAPS uses DAPS-1k for linear tasks and DAPS-4k for nonlinear tasks by default, and we retain these defaults (including hyperparameters) in our comparisons. Under these settings, DDiff attains higher reconstruction quality (PSNR/LPIPS; see Table 1) while requiring roughly half the sampling time on linear tasks and about one-sixth the sampling time on nonlinear tasks (see Fig. 5).
Appendix J More Qualitative Results
We provide additional qualitative results in Fig. 6 to 13.
Figure 6: Visual comparison of DDiff and baselines on super resolution task.
Figure 7: Visual comparison of DDiff and baselines on box inpainting task.
Figure 8: Visual comparison of DDiff and baselines on random inpainting task.
Figure 9: Visual comparison of DDiff and baselines on Gaussian deblurring task.
Figure 10: Visual comparison of DDiff and baselines on motion deblurring task.
Figure 11: Visual comparison of DDiff and baselines on phase retrieval task.
Figure 12: Visual comparison of DDiff and baselines on nonlinear deblurring task.
Figure 13: Visual comparison of DDiff and baselines on HDR task.